OpenStudy (anonymous):

A block of mass 0.245 kg is placed on top of a light, vertical spring of force constant 5 200 N/m and pushed downward so that the spring is compressed by 0.101 m. After the block is released from rest, it travels upward and then leaves the spring. To what maximum height above the point of release does it rise?

OpenStudy (anonymous):

Can you tell me what you've tried so far?

OpenStudy (anonymous):

Ok, I think you're going to have to use work and engergy to solve this. You can calculate the work done on the block by the spring (W=-1/2kx^2). Work is also equal to the change in kinetic engergy, so you can get the kinetic energy of the block immediately after it is released. Energy is conserved, so the final potential energy should be equal to the initial kinetic energy, and that will tell you how high the block can go. KE=1/2mv^2 and PE=mgh. Let me know if that makes sense, or if you need anything else.

OpenStudy (anonymous):

i tried calculating the work done by the spring on the block and getting Delta Kinetic energy. But the answer i get is wrong.

OpenStudy (anonymous):

What did you get for change in kinetic energy?